OpenStudy (anonymous):

Saxon wants to take a trip to Ireland. She plans to deposit $510 at the beginning of each quarter into an investment with a 2.95% interest rate, compounded quarterly. How much will she have in the account after 8 years?

OpenStudy (anonymous):

= $645.2

OpenStudy (anonymous):

Equation from wikipedia. \[FV = PV(1+i)^n\] Where FV = Future value (the value you end up with) PV= Present Value (the value you have now) i = interest rate (be careful, this should be converted into decimal form) n = times compounded (1 means it's compounded once a year)

OpenStudy (anonymous):

i got 4201.70 is tat correct?

OpenStudy (anonymous):

wait... I used the wrong formula. It didnt take the number of years into account.

OpenStudy (anonymous):

The formula for compound quarterly in 8 yrs: A = P (1+ r/ 4 ) ^ 32

OpenStudy (anonymous):

Just plug all numbers in!

OpenStudy (anonymous):

Thanks Chlorophyll. Kirihara, http://en.wikipedia.org/wiki/Compound_interest Section 2.2 on the wikipedia page explains the formula used.

OpenStudy (anonymous):

thanks guys !